Luebke Inc. has provided the following data for the month of November. The balance in the Finished Goods inventory account at the beginning of the month was $58,000 and at the end of the month was $30,600. The cost of goods manufactured for the month was $215,000. The actual manufacturing overhead cost incurred was $56,800 and the manufacturing overhead cost applied to Work in Process was $60,400. The company closes out any underapplied or overapplied manufacturing overhead to cost of goods sold. The adjusted cost of goods sold that would appear on the income statement for November is:

a) $238,800

b) 187,600

c) 242,400

d) 215,000
Answer #1

Ans is $238,800

Working:-

Cost of goods sold = opening balance of finished goods+ cost of goods manufactured - closing stock of finished goods
= 58000+215000-30600
= 242400
Actual manufacturing overheads = 56800
Overheads applied = 60400
Overheads overapplied = 60400-56800
= 3600
Adjusted cost of goods sold = cost of goods sold - overheads overapplied
= 242400-3600
= 238800
Hi mate,
I would be grateful to you if you can provide a thumbs up and write one beautiful comment. It will improve my rating and let me continue my journey here.
In case of doubt, please comment. I will consider myself fortunate if I can help you.
All the best for your bright future.